# Learning Robotic Policy with Imagined Transition: Mitigating the Trade-off between Robustness and Optimality

### v1

@@ VI Conclusion
In this paper, We propose LIT, a two-stage framework that mitigates the trade-off between robustness and optimality in RL-based locomotion. By introducing an imagined transition to policy learning, LIT enables policies trained with domain randomization to be guided by the desired gait in the simulator. Extensive experiments demonstrate significant improvements in velocity tracking under both simulators and real-world environments. LIT provides a new training paradigm when a large amount of domain randomness needs to be used for robustness. In future work, we will continue to investigate robotic scenarios exhibiting inherent trade-offs between precision and robustness, while exploring generalized implementations of this paradigm across broader application domains.

### v2

@@ VI Conclusion
In this paper, we propose LIT, a two-stage framework that aims to mitigate the trade-off between robustness and optimality in RL-based locomotion. By introducing an imagined transition to policy learning, LIT enables policies trained with domain randomization to be guided by the desired gait in the simulator. Extensive experiments demonstrate significant improvements in velocity tracking under both simulators and real-world environments. Inspired by reference signals in classical control methods, LIT provides the policy with reference actions and imagined next observations, supplies richer contextual information, enabling it to anticipate better and adapt to dynamic variations introduced by domain randomization. In future work, we will continue to investigate robotic scenarios exhibiting inherent trade-offs between precision and robustness, while exploring generalized implementations of this paradigm across broader application domains.